The gradient of the tangent to a curve is given by dy/dx = x2+1 / y2, y ≠0. The point P(-1,1) lies on the curve .

a)Find the equationof tangent to the cuve at P.

b) Solve the differential equation to find the equation of the curve in the form y= f(x).


1
Expert's answer
2021-09-20T16:04:59-0400

a) Let us find the equation "y=y(x_0)+y'(x_0)(x-x_0)" of tangent to the curve at "P(x_0,y_0)," where "x_0=-1,\ y_0=1." Since "y'(1)=\frac{(-1)^2+1}{1^2}=2," it follows that the equation of tangent is "y=1+2(x+1)" or "y=2x+3."


b) Let us solve the differential equation "\frac{dy}{dx} = \frac{x^2+1}{ y^2}," which is equivalent to "y^2dy=(x^2+1)dx". It follows that "\int y^2dy=\int(x^2+1)dx," and hence "\frac{y^3}3=\frac{x^3}{3}+x+C." We conclude that "y^3=x^3+3x+C_1," and therefore, the general solution of the differential equation is

"y=\sqrt[3]{x^3+3x+C_1}."